OpenStudy (anonymous):

g=[(-7,-5) , (-4,-7) , (5,0) , (9,6)] h(x)=2x-3 find g^-1(-7)=? h^-1(x)=? (h times h^-1)(3)=?

OpenStudy (anonymous):

\[g(-4)=-7\] so \[g^{-1}(-7)=-4\]

OpenStudy (anonymous):

h says mulitiply by 2, subtract 3 so \[h^{-1}\] will say add 3 divide by 2 \[h^{-1}(x)=\frac{x+3}{2}\]

OpenStudy (anonymous):

if that is too annoying think about solving this 2x+3=11 steps are subtract 3 divide by 2 to get x = 4 now if we have y = 2x +3 and we want to solve for x we write y - 3 = 2x or x = (y-3)/2 and that is your inverse

OpenStudy (anonymous):

and finally it is not times, it is "of" not \[h\times h^{-1}(x)\] but rather \[h(h^{-1}(x))\] and \[h(h^{-1}(x))=x\] because when you compose a function with its inverse you get what you started with